# 3D Shape Generation: A Survey

### v1

@@ 4.1 Feedforward Generation
Feedforward generation refers to the direct synthesis of 3D representations through generative models [LYC+25]. In this section, we explore these methods based on their generative models as shown in FIG 4, which include (a) Generative Adversarial Networks (GANs), (b) Variational Autoencoders (VAEs), (c) Normalizing Flows, (d) Autoregressive Models, and (e) Diffusion Models.
Figure 4: Exemplary pipelines of feedforward 3D generation methods. Source [LZK+24].
@@ 4.1.1 Generative Adversarial Networks (GANs)
Generative Adversarial Networks (GANs) [GPM+14] have demonstrated remarkable outcomes in the domain of 3D shape generation, consisting of a generator $G(\cdot)$ and a discriminator $D(\cdot)$, where the discriminator tries to classify real objects and objects synthesized by the generator, and the generator attempts to fool the discriminator [WZX+16].
The core idea is to integrate GANs with 3D shape representations (as discussed in Section 3), enabling the generation of geometries such as point clouds (l-GAN/r-GAN [ADMG18], tree-GAN [SPK19], SP-GAN [LLHF21]), voxel grids (3D-GAN [WZX+16], Z-GAN [KKR19]), meshes (MeshGAN [CBZ+19]), or SDF (SurfGen [LLZL21], SDF-StyleGan [ZLWT22]). In this context, GANs serve as the underlying generative framework that learns to synthesize 3D shapes by mapping latent codes to complex geometric structures [LLHF21].
In the context of 3D shape generation, prior GAN-based methods such as 3D-GAN [WZX+16], l-GAN/r-GAN [ADMG18], and Multi-chart Generation [BMK+18] leverage explicit representations of real data to guide the training of generator networks. These approaches employ discriminators that assess 3D structures to supervise the generator, encouraging the synthesis of shapes that resemble the geometric realism of the training data. While these methods are capable of generating high-resolution shapes, they still struggle with thin structures and implausible objects. To address this limitation, SurfGen [LLZL21] was introduced as an extension of implicit GANs with differentiable surface-aware rendering. SurfGen adopts DeepSDF [PFS+19] as the generator within its proposed framework. To enable surface-level discrimination on objects represented as SDFs, the zero-level isosurface is first identified. Then, the marching cubes algorithm [LC87] is applied to extract an explicit triangle mesh from the continuous SDF representation. For differentiable isosurface extraction, SurfGen leverages the MeshSDF method introduced in [RLR+20].
@@ 4.1.2 Variational Autoencoders (VAEs)
Suppose we have a random variable $X$ for which we are building generative models. The variational autoencoder (VAE) is a framework that allows one to learn $P(X)$ from a dataset of observations of $X$ [KW14, RMW14, YHH+19]. The VAE models the data distribution via a latent variable $z$ with a prior distribution $P_{\psi}(z)$, and a decoder $P_{\theta}(X\,|\,z)$ which captures the (hopefully simpler) distribution of $X$ given $z$. During training, it additionally learns an inference model (or encoder) $Q_{\phi}(z\,|\,X)$. The encoder and decoder are jointly trained to maximize a lower bound on the log-likelihood of the observations
@@
In this context, VAEs [KW14] are probabilistic generative models composed of two neural networks: an encoder and a decoder. The encoder maps an input data point to a latent space by producing the parameters of a variational distribution–tipycally a multivariate Gaussian–from which latent vectors can be sampled. This enables the encoder to generate multiple samples that originate from the same underlying distribution. The decoder then maps these latent variables back to the data space, reconstructing or generating new samples. Both networks are typically trained jointly using the reparametrization trick, which allows for gradient-based optimization through the sampling process. Although the noise model’s variance can also be learned, it is sometimes treated separately for stability or modeling purposes [LZK+24].
@@
[BLRW16] introduced a voxel-based VAE alongside a graphical user interface for exploring the latent space of 3D generative models. Their framework also includes a deep voxel-based convolutional neural network for classification. In parallel, SDM-Net [GYW+19] focuses on generating structured meshes composed of deformable parts. The method uses one VAE network to model parts and another to model the whole object. Building upon this idea, TM-Net [GWY+21] extends the approach to generate texture maps for meshes in a part-aware manner.
@@
In the context of 3D shape generation, GRASS [LXC+17] proposes a structure-aware generation framework, which learns a latent space of shape layouts (parts and their connections) using a recursive decoder that splits shape parts hierarchically. GRASS was the first to generate 3D shapes as a composition of parts, producing plausible novel combinations (e.g., new chair designs) and enabling interpolation at the part-structure level. Subsequently, ShapeVAE [NW17] was one of the first VAE-based models proposed for 3D shape generation. It introduced a probabilistic latent space over 3D object geometry, enabling the generation of novel shape variants through latent sampling. However, a key limitation of the model was its low output resolution. ShapeVAE also incorporated part segmentation into its latent representation, allowing for intuitive shape analogies and part-level manipulations. Additional VAE-based models have since been proposed, including AtlasNet [GFK+18], which reconstructs surfaces from learned parametrizations; SetVAE [KYLH21], which extends the VAE framework to handle unordered point sets using hierarchical structure-aware encoding; and Michelangelo [ZLC+23], which leverages part-aware priors in conditional generative tasks.
@@
In synthesis, VAEs are often used in combination with other generative paradigms to enhance structure modeling and multimodal conditioning. Compared to GAN-based approaches, the training of VAEs is considerably more stable than GANs. However, VAEs tend to produce more blurred results compared to GANs [LZK+24].
@@ 4.1.3 Normalizing Flows
A normalizing flow [RM15] is a series of invertible mappings that transform an initial known distribution $p_{Z}(z)$ to a complex data distribution $p_{X}(x)$ [KLK+20].
In the context of 3D shape generation, PointFlow [YHH+19] learns the distribution of shapes, each shape itself being a distribution of points using continuous normalizing flows [CRBD18, GCB+18]. PointFlow was the first to show that flow-based models can generate 3D point sets with improved coverage (diversity), though at a higher computational cost. To address this limitation, the Discrete Point Flow (DPF) networks [KBV20] were introduced, which use affine coupling layers (a discrete flow) to map noise to point cloud coordinates. DPF networks achieve a significant speed-up in generation with comparable fidelity, making flow models more practical for 3D.
SoftFlow [KLK+20] is a framework for training normalizing flows on the manifold. In order to bridge the gap between the dimensions of the data and the target latent variable, the authors propose a method to estimate a conditional distribution of the perturbed data. SoftFlow resolves the dimension mismatch and improves the stability of flow-based 3D shape generation.
Subsequent works–such as Point Straight Flow (PSF) [WWG+23] and EAGLE [WZYW25]–have further advanced the use of normalizing flows for 3D shape generation. In particular, EAGLE proposes adaptive continuous normalizing flows by introducing a control mechanism with adaptive learning rates to the bias term.
@@ 4.1.4 Autoregressive Models
Flow-based models and auto-regressive models can both perform exact likelihood evaluation, while flow-based models are much more efficient to sample from [YHH+19].
Autoregressive networks model current values as a function of their own previous values [SWL+20]. According to [LZK+24], a 3D object can be represented as a joint probability of the occurrences of multiple 3D elements:
@@
In the context of 3D shape generation, several studies have explored autoregressive modeling as a generative approach. PointGrow [SWL+20], in turn, employs a Transformer-like self-attention mechanism to autoregressively predict points step by step, each conditioned on the previously generated points. It captures long-range dependencies and yields diverse shapes, but is inherently slower to sample. Inspired by the autoregressive design of PointGrow, PolyGen [NGEB20] introduces a transformer-based framework for generating 3D meshes. It employs two sequential transformer networks: a vertex model, which unconditionally models mesh vertices, and a face model, which models the mesh faces conditioned on input vertices. This approach enables the structured synthesis of meshes while maintaining geometric consistency through autoregressive modeling.
Subsequent works have extended autoregressive modeling to various 3D representations. AutoSDF [MCST22] generates 3D shapes represented by volumetric truncated-signed distance functions (T-SDF). It employs a VQ-VAE framework to learn a quantized codebook that captures local geometric patterns within T-SDFs. Shapes are encoded as discrete tokens from this codebook and modeled using a transformer-based architecture in a non-sequential autoregressive manner. Also, AutoSDF is capable of completing shapes based on images or text.
Additional approaches include GraphRNN [YYR+18], ScanComplete [DRB+18], StructureNet [MGY+19], ShapeFormer [YLM+22], ShapeCrafter [FZC+23], Octree Transformer [IKK23], and the more recent PointARU [MWZ+25], which adopts a coarse-to-fine autoregressive up-sampling approach, eliminating the need to define a fixed generation order for point sequences. These methods explore diverse input structures–such as graphs, hierarchical trees, and octrees–while leveraging autoregressive priors to generate semantically and geometrically coherent shapes.
@@ 4.1.5 Diffusion Models
Diffusion models [SWMG15, HJA20] are a class of generative models that convert simple known distributions (e.g., a Gaussian) into complex data distributions. These models generate data by simulating a reverse diffusion process, gradually transforming a noise sample into a structured output. Their application to 3D shape generation has shown promising performance in capturing fine geometric details and enabling flexible conditioning mechanisms [RBL+22, LZK+24].
@@
Several methods have explored diffusion directly in the point cloud space. For instance, ShapeGF [CYA+20] builds upon a denoising score-matching framework to learn the underlying data distributions of point clouds. PVD [ZDW21] introduces a bidirectional diffusion model that learns mappings between point clouds and voxel grids, enabling both generation and reconstruction tasks. Similarly, DPM [LH21b] learns a noise-conditioned score function to refine noisy point clouds into clean samples progressively. Building on these foundations, LION [ZVW+22] introduces a latent diffusion model specifically tailored for point cloud generation. By operating in a learned latent space, LION improves sampling efficiency and scalability while maintaining geometric fidelity. TIGER [RKLL24] further advances this direction by incorporating a time-varying denoising process in a single triangle-grid representation, achieving high-resolution shape synthesis with fine topological structure.
@@
Beyond point clouds, MeshDiffusion [LFB+22] and SLIDE [LWA+23] employ score-based and controllable diffusion strategies, respectively, to generate meshes directly with controllable topology and attributes. Tetrahedral Diffusion Models [KPWS25] explore volumetric mesh generation using tetrahedral elements, enabling continuous and differentiable surface extraction.
@@
Additional works have extended diffusion to hybrid and latent representations. Shap-E [JN23] learns a unified latent field for conditional generation from images to text. Wavelet Diffusion [HHL+24a] and WaLa [SKR+24] introduce wavelet-based latent diffusion to improve resolution and multi-scale control. SALAD [KYNS23] focuses on part-level generation using latent diffusion models that can synthesize individual object components. Point-E [NJD+22] combines the power of GLIDE [NDR+22] for conditional image generation with a diffusion model-based decoder that produces point clouds from synthesized views. XCube [RHZ+24] introduces a scalable approach for 3D shape generation based on sparse voxel hierarchies, enabling memory-efficient synthesis of high-resolution geometry across large and diverse object categories.
@@
Recent advances also include models that further expand the capabilities of latent diffusion for 3D shape generation. GPLD3D [DZG+24] proposes latent diffusion with enforced geometric and physical priors to improve the realism and plausibility of generated shapes. The triplane-based 3D-aware Diffusion model with TransFormer (DiffTF) [CHW+23] introduces a transformer-based latent diffusion framework capable of generating shapes across a wide range of object categories. Sin3DM [WLVZ23] focuses on single-shot learning, training a diffusion model from just a single 3D textured shape. Finally, Direct3D [WLZ+24] presents a scalable image-to-3D generation approach based on a 3D latent diffusion transformer, enabling efficient and high-quality synthesis conditioned on images.
@@
Together, these approaches showcase the versatility and scalability of diffusion models for handling diverse 3D shape representations, including explicit, implicit, and hybrid representations. As research in this area continues to advance, it is expected that diffusion models will play a crucial role in pushing the boundaries of 3D shape generation across a wide range of applications [LZK+24].
@@ 5 Open-Source Datasets for 3D Generation
To address fine-pose estimation of objects in 2D images given a 3D model, the dataset introduced in  [LPT13], named as IKEA objects, includes 800 images and 225 IKEA-style CAD models sourced from Google Warehouse. Other datasets have been introduced for specific applications, including robotic manipulation [CWS+15, MCL20], 3D shape retrieval (Shrec’14 [GL14]), single-image 3D shape reconstruction (Pix3D [SWZ+18]), and scanned objects (Google Scanned Objects [DFK+22]). The BigBird dataset [SSN+14] contributes a large-scale collection of 3D object instances along with multi-view RGB images, depth maps, camera pose annotations, and object segmentations.
@@
One of the most widely used datasets in 3D generation is ShapeNet [CFG+15], which includes two main subsets: ShapeNetCore and ShapeNetSem. The full dataset comprises nearly 3 million textured 3D computer-aided design (CAD) models, spanning 3,135 object categories. ShapeNetCore, the most commonly used subset, contains approximately 51,300 models across 55 object categories, each with manually verified annotations for category labels and object alignment. The standard experimental setup for evaluating generative models on ShapeNet is often adapted from the benchmark introduced by PointFlow.
ModelNet [ZSK+15] is another well-known, extensive shape dataset of 3D CAD models. It is one of the most widely used benchmarks for point cloud analysis. It contains 151,128 objects belonging to 660 categories.
ObjectNet3D [XKC+16] contains 3D shapes and 2D images. The 2D images are aligned with the corresponding 3D shapes. ObjectNet consists of 100 categories, 90,127 images, 201,88 objects in these images, and 44,147 3D shapes. The objects in the 2D images are aligned with the 3D shapes, and the alignment provides both accurate 3D pose annotation and the closest 3D shape annotation for each 2D object.
@@
PartNet [MZC+19] is a consistent, large-scale dataset on top of ShapeNet, with fine-grained, hierarchical, instance-level 3D part information. PartNet comprises 573,585 part instances over 26,671 3D models covering 24 object categories. It is commonly used for point cloud-related tasks.
@@
3D-FUTURE [FJG+21] is a large-scale dataset that contains 20,240 realistic indoor images and the associated 9,992 unique 3D furniture models with rich geometry details and high-resolution textures. 3D-FUTURE also offers instance segmentation annotation and image rendering information, including six degrees of freedom (6DoF).
The Amazon Berkeley Objects (ABO) [CGD+22] dataset was designed to help bridge the gap between real and virtual 3D worlds. ABO is derived from Amazon.com product listings. It contains 147,702 product listings associated with 398,212 unique catalog images. ABO also includes 7,953 products with corresponding artist-designed 3D meshes.
Objaverse [DSS+23] is a large-scale, high-quality dataset of 3D objects with rich annotations, including natural language captions. Objaverse contains over 800K textured 3D models spanning a wide variety of categories and use cases. Its successor, Objaverse-XL [DLW+23] further extends Objaverse to a larger 3D dataset of 10.2M 3D objects, establishing it as one of the largest publicly available corpora for multimodal 3D learning.
These large-scale 3D datasets have the potential to facilitate large-scale training and boost the performance of 3D generation [LZK+24], specifically in the context of 3D shape generation.

### v2

@@ 4.1 Feedforward Generation
Feedforward generation refers to the direct synthesis of 3D representations through generative models [LYC25]. In this section, we explore these methods based on their generative models as shown in Figure 3, which include (a) Generative Adversarial Networks (GANs), (b) Variational Autoencoders (VAEs), (c) Normalizing Flows, (d) Autoregressive Models, and (e) Diffusion Models.
Figure 3: Exemplary pipelines of feedforward 3D generation methods.
@@ 4.1.1 Generative Adversarial Networks (GANs)
Generative Adversarial Networks (GANs) [GPM14] have demonstrated remarkable outcomes in the domain of 3D shape generation, consisting of a generator $G(\cdot)$ and a discriminator $D(\cdot)$, where the discriminator tries to classify real objects and objects synthesized by the generator, and the generator attempts to fool the discriminator [WZX16].
The core idea is to integrate GANs with 3D shape representations (as discussed in Section 3), enabling the generation of geometries such as point clouds (l-GAN/r-GAN [ADMG18], tree-GAN [SPK19], SP-GAN [LLHF21]), voxel grids (3D-GAN [WZX16], Z-GAN [KKR19]), meshes (MeshGAN [CBZ19]), or SDF (SurfGen [LLZL21], SDF-StyleGan [ZLWT22]). In this context, GANs serve as the underlying generative framework that learns to synthesize 3D shapes by mapping latent codes to complex geometric structures [LLHF21].
In the context of 3D shape generation, prior GAN-based methods such as 3D-GAN [WZX16], l-GAN/r-GAN [ADMG18], and Multi-chart Generation [BMK18] leverage explicit representations of real data to guide the training of generator networks. These approaches employ discriminators that assess 3D structures to supervise the generator, encouraging the synthesis of shapes that resemble the geometric realism of the training data. While these methods are capable of generating high-resolution shapes, they still struggle with thin structures and implausible objects. To address this limitation, SurfGen [LLZL21] was introduced as an extension of implicit GANs with differentiable surface-aware rendering. SurfGen adopts DeepSDF [PFS19] as the generator within its proposed framework. To enable surface-level discrimination on objects represented as SDFs, the zero-level isosurface is first identified. Then, the marching cubes algorithm [LC87] is applied to extract an explicit triangle mesh from the continuous SDF representation. For differentiable isosurface extraction, SurfGen leverages the MeshSDF method introduced in [RLR20].
@@ 4.1.2 Variational Autoencoders (VAEs)
Suppose we have a random variable $X$ for which we are building generative models. The variational autoencoder (VAE) is a framework that allows one to learn $P(X)$ from a dataset of observations of $X$ [KW14, RMW14, YHH19]. The VAE models the data distribution via a latent variable $z$ with a prior distribution $P_{\psi}(z)$, and a decoder $P_{\theta}(X\,|\,z)$ which captures the (hopefully simpler) distribution of $X$ given $z$. During training, it additionally learns an inference model (or encoder) $Q_{\phi}(z\,|\,X)$. The encoder and decoder are jointly trained to maximize a lower bound on the log-likelihood of the observations
@@
In this context, VAEs [KW14] are probabilistic generative models composed of two neural networks: an encoder and a decoder. The encoder maps an input data point to a latent space by producing the parameters of a variational distribution–tipycally a multivariate Gaussian–from which latent vectors can be sampled. This enables the encoder to generate multiple samples that originate from the same underlying distribution. The decoder then maps these latent variables back to the data space, reconstructing or generating new samples. Both networks are typically trained jointly using the reparametrization trick, which allows for gradient-based optimization through the sampling process. Although the noise model’s variance can also be learned, it is sometimes treated separately for stability or modeling purposes [LZK24].
@@
[BLRW16] introduced a voxel-based VAE alongside a graphical user interface for exploring the latent space of 3D generative models. Their framework also includes a deep voxel-based convolutional neural network for classification. In parallel, SDM-Net [GYW19] focuses on generating structured meshes composed of deformable parts. The method uses one VAE network to model parts and another to model the whole object. Building upon this idea, TM-Net [GWY21] extends the approach to generate texture maps for meshes in a part-aware manner.
@@
In the context of 3D shape generation, GRASS [LXC17] proposes a structure-aware generation framework, which learns a latent space of shape layouts (parts and their connections) using a recursive decoder that splits shape parts hierarchically. GRASS was the first to generate 3D shapes as a composition of parts, producing plausible novel combinations (e.g., new chair designs) and enabling interpolation at the part-structure level. Subsequently, ShapeVAE [NW17] was one of the first VAE-based models proposed for 3D shape generation. It introduced a probabilistic latent space over 3D object geometry, enabling the generation of novel shape variants through latent sampling. However, a key limitation of the model was its low output resolution. ShapeVAE also incorporated part segmentation into its latent representation, allowing for intuitive shape analogies and part-level manipulations.
Additional VAE-based models have since been proposed, including AtlasNet [GFK18], which reconstructs surfaces from learned parametrizations; SetVAE [KYLH21], which extends the VAE framework to handle unordered point sets using hierarchical structure-aware encoding; Michelangelo [ZLC23], which leverages part-aware priors in conditional generative tasks; TAR3D [ZLL25], which extends the learning capabilities of the next-token prediction paradigm to conditional 3D object generation, enabling more effective sequence-based modeling of 3D structures; and COD-VAE [CYJK25], which introduces a representation framework that encondes 3D shapes into a set of compact 1D latent vectors, providing a lightweight yet expressive latent space for generation and reconstruction.
@@
In synthesis, VAEs are often used in combination with other generative paradigms to enhance structure modeling and multimodal conditioning. Compared to GAN-based approaches, the training of VAEs is considerably more stable than GANs. However, VAEs tend to produce more blurred results compared to GANs [LZK24].
@@ 4.1.3 Normalizing Flows
A normalizing flow [RM15] is a series of invertible mappings that transform an initial known distribution $p_{Z}(z)$ to a complex data distribution $p_{X}(x)$ [KLK20].
In the context of 3D shape generation, PointFlow [YHH19] learns the distribution of shapes, each shape itself being a distribution of points using continuous normalizing flows [CRBD18, GCB18]. PointFlow was the first to show that flow-based models can generate 3D point sets with improved coverage (diversity), though at a higher computational cost. To address this limitation, the Discrete Point Flow (DPF) networks [KBV20] were introduced, which use affine coupling layers (a discrete flow) to map noise to point cloud coordinates. DPF networks achieve a significant speed-up in generation with comparable fidelity, making flow models more practical for 3D.
SoftFlow [KLK20] is a framework for training normalizing flows on the manifold. In order to bridge the gap between the dimensions of the data and the target latent variable, the authors propose a method to estimate a conditional distribution of the perturbed data. SoftFlow resolves the dimension mismatch and improves the stability of flow-based 3D shape generation.
Subsequent works–such as Point Straight Flow (PSF) [WWG23] and EAGLE [WZYW25]–have further advanced the use of normalizing flows for 3D shape generation. In particular, EAGLE proposes adaptive continuous normalizing flows by introducing a control mechanism with adaptive learning rates to the bias term.
@@ 4.1.4 Autoregressive Models
Flow-based models and auto-regressive models can both perform exact likelihood evaluation, while flow-based models are much more efficient to sample from [YHH19].
Autoregressive networks model current values as a function of their own previous values [SWL20]. According to [LZK24], a 3D object can be represented as a joint probability of the occurrences of multiple 3D elements:
@@
In the context of 3D shape generation, several studies have explored autoregressive modeling as a generative approach. PointGrow [SWL20], in turn, employs a Transformer-like self-attention mechanism to autoregressively predict points step by step, each conditioned on the previously generated points. It captures long-range dependencies and yields diverse shapes, but is inherently slower to sample. Inspired by the autoregressive design of PointGrow, PolyGen [NGEB20] introduces a transformer-based framework for generating 3D meshes. It employs two sequential transformer networks: a vertex model, which unconditionally models mesh vertices, and a face model, which models the mesh faces conditioned on input vertices. This approach enables the structured synthesis of meshes while maintaining geometric consistency through autoregressive modeling.
Subsequent works have extended autoregressive modeling to various 3D representations. AutoSDF [MCST22] generates 3D shapes represented by volumetric truncated-signed distance functions (T-SDF). It employs a VQ-VAE [vVk17] framework to learn a quantized codebook that captures local geometric patterns within T-SDFs. Shapes are encoded as discrete tokens from this codebook and modeled using a transformer-based architecture in a non-sequential autoregressive manner. Also, AutoSDF is capable of completing shapes based on images or text.
Additional approaches include GraphRNN [YYR18], ScanComplete [DRB18], StructureNet [MGY19], ShapeFormer [YLM22], ShapeCrafter [FZC23], Octree Transformer [IKK23], and PointARU [MWZ25], which adopts a coarse-to-fine autoregressive up-sampling approach, eliminating the need to define a fixed generation order for point sequences. In addition, ShapeGPT [YCZ25] introduces a multi-modal framework that leverages large pre-trained language models to address multiple tasks, including text-to-shape, shape-to-text, shape completion, and shape editing.
These methods explore diverse input structures–such as graphs, hierarchical trees, and octrees–while leveraging autoregressive priors to generate semantically and geometrically coherent shapes.
@@ 4.1.5 Diffusion Models
Diffusion models [SWMG15, HJA20] are a class of generative models that convert simple known distributions (e.g., a Gaussian) into complex data distributions. These models generate data by simulating a reverse diffusion process, gradually transforming a noise sample into a structured output. Their application to 3D shape generation has shown promising performance in capturing fine geometric details and enabling flexible conditioning mechanisms [RBL22, LZK24].
@@
Several methods have explored diffusion directly in the point cloud space. For instance, ShapeGF [CYA20] builds upon a denoising score-matching framework to learn the underlying data distributions of point clouds. PVD [ZDW21] introduces a bidirectional diffusion model that learns mappings between point clouds and voxel grids, enabling both generation and reconstruction tasks. Similarly, DPM [LH21b] learns a noise-conditioned score function to refine noisy point clouds into clean samples progressively. Building on these foundations, LION [ZVW22] introduces a latent diffusion model specifically tailored for point cloud generation. By operating in a learned latent space, LION improves sampling efficiency and scalability while maintaining geometric fidelity. TIGER [RKLL24] further advances this direction by incorporating a time-varying denoising process in a single triangle-grid representation, achieving high-resolution shape synthesis with fine topological structure.
@@
Beyond point clouds, MeshDiffusion [LFB22] and SLIDE [LWA23] employ score-based and controllable diffusion strategies, respectively, to generate meshes directly with controllable topology and attributes. Tetrahedral Diffusion Models [KPWS25] explore volumetric mesh generation using tetrahedral elements, enabling continuous and differentiable surface extraction.
@@
Additional works have extended diffusion to hybrid and latent representations. Shap-E [JN23] learns a unified latent field for conditional generation from images to text. Wavelet Diffusion [HHL24a] and WaLa [SKR24] introduce wavelet-based latent diffusion to improve resolution and multi-scale control. SALAD [KYNS23] and CoPart [DDC25] focus on part-level generation using latent diffusion models that can synthesize individual object components. Point-E [NJD22] combines the power of GLIDE [NDR22] for conditional image generation with a diffusion model-based decoder that produces point clouds from synthesized views. XCube [RHZ24] introduces a scalable approach for 3D shape generation based on sparse voxel hierarchies, enabling memory-efficient synthesis of high-resolution geometry across large and diverse object categories.
@@
Recent advances also include models that further expand the capabilities of latent diffusion for 3D shape generation. GPLD3D [DZG24] proposes latent diffusion with enforced geometric and physical priors to improve the realism and plausibility of generated shapes. The triplane-based 3D-aware Diffusion model with TransFormer (DiffTF) [CHW23] introduces a transformer-based latent diffusion framework capable of generating shapes across a wide range of object categories. Sin3DM [WLVZ23] focuses on single-shot learning, training a diffusion model from just a single 3D textured shape. Finally, Direct3D [WLZ24] presents a scalable image-to-3D generation approach based on a 3D latent diffusion transformer, enabling efficient and high-quality synthesis conditioned on images. Likewise, Hi3DGen [YWL25] introduces a normal bridging strategy that further enhances fidelity and efficiency in image-conditioned 3D generation.
@@
Together, these approaches showcase the versatility and scalability of diffusion models for handling diverse 3D shape representations, including explicit, implicit, and hybrid representations. As research in this area continues to advance, it is expected that diffusion models will play a crucial role in pushing the boundaries of 3D shape generation across a wide range of applications [LZK24].
@@ 5 Open-Source Datasets for 3D Generation
To address fine-pose estimation of objects in 2D images given a 3D model, the dataset introduced in  [LPT13], named as IKEA objects, includes 800 images and 225 IKEA-style CAD models sourced from Google Warehouse. Other datasets have been introduced for specific applications, including robotic manipulation [CWS15, MCL20], 3D shape retrieval (Shrec’14 [GL14]), single-image 3D shape reconstruction (Pix3D [SWZ18]), and scanned objects (Google Scanned Objects [DFK22]). The BigBird dataset [SSN14] contributes a large-scale collection of 3D object instances along with multi-view RGB images, depth maps, camera pose annotations, and object segmentations.
@@
One of the most widely used datasets in 3D generation is ShapeNet [CFG15], which includes two main subsets: ShapeNetCore and ShapeNetSem. The full dataset comprises nearly 3 million textured 3D computer-aided design (CAD) models, spanning 3,135 object categories. ShapeNetCore, the most commonly used subset, contains approximately 51,300 models across 55 object categories, each with manually verified annotations for category labels and object alignment. The standard experimental setup for evaluating generative models on ShapeNet is often adapted from the benchmark introduced by PointFlow.
ModelNet [ZSK15] is another well-known, extensive shape dataset of 3D CAD models. It is one of the most widely used benchmarks for point cloud analysis. It contains 151,128 objects belonging to 660 categories.
ObjectNet3D [XKC16] contains 3D shapes and 2D images. The 2D images are aligned with the corresponding 3D shapes. ObjectNet consists of 100 categories, 90,127 images, 201,88 objects in these images, and 44,147 3D shapes. The objects in the 2D images are aligned with the 3D shapes, and the alignment provides both accurate 3D pose annotation and the closest 3D shape annotation for each 2D object.
@@
PartNet [MZC19] is a consistent, large-scale dataset on top of ShapeNet, with fine-grained, hierarchical, instance-level 3D part information. PartNet comprises 573,585 part instances over 26,671 3D models covering 24 object categories. It is commonly used for point cloud-related tasks.
@@
3D-FUTURE [FJG21] is a large-scale dataset that contains 20,240 realistic indoor images and the associated 9,992 unique 3D furniture models with rich geometry details and high-resolution textures. 3D-FUTURE also offers instance segmentation annotation and image rendering information, including six degrees of freedom (6DoF).
The Amazon Berkeley Objects (ABO) [CGD22] dataset was designed to help bridge the gap between real and virtual 3D worlds. ABO is derived from Amazon.com product listings. It contains 147,702 product listings associated with 398,212 unique catalog images. ABO also includes 7,953 products with corresponding artist-designed 3D meshes.
Objaverse [DSS23] is a large-scale, high-quality dataset of 3D objects with rich annotations, including natural language captions. Objaverse contains over 800K textured 3D models spanning a wide variety of categories and use cases. Its successor, Objaverse-XL [DLW23] further extends Objaverse to a larger 3D dataset of 10.2M 3D objects, establishing it as one of the largest publicly available corpora for multimodal 3D learning.
More recently, [DDC25] introduces PartVerse, a large-scale 3D part dataset comprising 91k high-quality parts from 12k objects with detailed descriptions. PartVerse is a curated and annotated dataset from Objaverse, and exhibits enhanced diversity in 175 categories and includes realistic textures, making it a valuable benchmark for part-level 3D understanding and generation.
[…]
These large-scale 3D datasets have the potential to facilitate large-scale training and boost the performance of 3D generation [LZK24], specifically in the context of 3D shape generation.